let Tr=r th term of an ap. m.Tm=n.Tn,show that Tm+n=0

Grade:11

1 Answers

Vikas TU
5 years ago
r th term of an ap= Tr 
m th term of an ap= Tm 
n th term of an ap= Tn 
As indicated by question m.Tm=n.Tn 
m(a + (m-1)d = n(a + (n-1)d) 
Where an is the principal term and d is the normal distinction 
mama – na + (m^2 – m)d - (n^2 –n)d 
a(m-n) + d(m-n){m+n-1} = 0 
a + (m+n-1)d = 0 
Tm+n=0 
Thus Proved